OpenStudy (haaileyyy):

Please Help Find the Product: (a^2-b)(5a^2+6b) (

OpenStudy (mariamikayla):

\[(a^2-b)(5a^2+6b)=a^2*5a^2+a^2*6b-b*5a^2-b*6b= \\\\ =5a^4+6ba^2-5a^2b-6b^2=\\\\=5a^4+a^2b-6b^2\]

OpenStudy (haaileyyy):

Thank You!